# The non-perturbative adiabatic invariant is all you need

### v1

@@ Abstract
[…]
Figure 1: Fusion-born $\alpha$-particle $\epsilon=\rho_{0}/L_{0}$ vs magnetic field strength $B_{0}$ in a device with scale length $L_{0}=1\,\text{m}$
Figure 2: (Top) Poincaré sections ($z=v_{z}=v_{y}=0,v_{x}>0$) of the full-orbit dynamics of Eq. (1) (black) and level sets of an adiabatic invariant truncation (color) for different values of $\epsilon$. Second-order truncation is used in (a); first-order truncation is used in (b) and (c). (Bottom) Relative error time traces of three truncations of adiabatic invariant series for the same $\epsilon$ values. Underlying trajectories are denoted by a blue cross in (a)-(c).
[…]
For particles with $\epsilon\in\{0.05,0.15,0.27\}$ (cf. Fig. 1) moving in the simple $z$-symmetric field geometry $\bm{B}=B(x,y)\,\bm{e}_{z}$, where
@@
[…]
the leading-order perturbative adiabatic invariant $\mu_{0}={m\,v_{\perp}^{2}}/{(2|\bm{B}|)}$ suffers strikingly-large variations in time. Worse yet, for $\epsilon\in\{0.15,0.27\}$, when $\mu_{0}$ is replaced with higher-order asymptotic corrections Burby_gc_2013 conservation need not improve, indicating collapse of the adiabatic invariant’s optimal truncation order Berry_1991 . See Fig. 2. It follows that the guiding center asymptotic expansion may grossly mischaracterize $\alpha$-particle trajectories in today’s stellarator optimization codes. On the other hand, the phase portraits in Fig. 2 reveal that a large fraction of $\alpha$ trajectories do enjoy an adiabatic invariant, even though it cannot be captured using standard guiding center theory; its level sets coincide with the invariant circles in the phase portrait. (Note there is probably no true smooth invariant, even for $\epsilon\in\{0.05,0.15\}$, due to presence of very thin chaotic bands.) This suggests there may yet be a good notion of adiabatic invariant, and perhaps even guiding center dynamics, for $\alpha$’s that break the traditional guiding center model. Whatever that notion may be, collapse of the optimal truncation order indicates it lies beyond the reach of traditional asymptotic expansions.
@@
In this Letter we describe a non-perturabative guiding center model suitable for $\alpha$-particles in stellarators. First we deduce non-perturbative guiding center equations of motion assuming the non-perturbative adiabatic invariant $\mathcal{J}$ is known. Remarkably, these equations are completely determined by first-order derivatives of $\mathcal{J}$ and the magnetic field. They enjoy a Hamiltonian structure comparable with that of the usual asymptotic theory Littlejohn_1981 . Then we describe a data-driven method for learning $\mathcal{J}$ from a dataset of full-orbit $\alpha$-particle trajectories. We apply this method to $\alpha$ dynamics in the fields underlying Fig. 2 and find that the non-perturbative guiding center model determined by our learned $\mathcal{J}$ significantly outperforms the standard guiding center expansion. Our work establishes the need for a non-perturbative adiabatic invariant that can be applied in any magnetic field configuration; the method of finding $\mathcal{J}$ reported here can only be applied on a per-magnetic-field basis. The fact that such a general-purpose adiabatic invariant exists in the perturbative regime suggests a non-perturbative general-purpose adiabatic invariant may exist as well.
@@
In a seminal paper, Kruskal Kruskal_1962 showed that the traditional guiding center expansion originates from a hidden perturbative $U(1)$-symmetry found in the ordinary differential equations describing single-particle motion, $X=\dot{\bm{x}}\cdot\partial_{\bm{x}}+\dot{\bm{v}}\cdot\partial_{\bm{v}}$, where $\dot{\bm{v}}=\bm{v}\times\bm{B}$ and $\dot{\bm{x}}=\epsilon\,\bm{v}$. He dubbed the infinitesimal generator of the perturbative hidden symmetry the roto rate $R$ and showed that its formal expansion in powers of $\epsilon$ is uniquely determined to all orders, with first term given by $R_{0}=\bm{v}\times\bm{b}\cdot\partial_{\bm{v}}$.
@@
i.e. $\{E,\mathcal{J}\}=0$. We will refer to $\mathcal{J}$ as the non-perturbative adiabatic invariant.
@@
To ensure topological similarity with the known $U(1)$-symmetry at $\epsilon=0$, we also assume (II) that each $U(1)$-orbit generated by $\mathcal{R}$ transversally intersects a certain phase space Poincaré section $\Sigma$ exactly once. To define $\Sigma$, introduce phase space coordinates $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ such that $\bm{v}=v_{\parallel}\,\bm{b}+v_{\perp}\,(\cos\zeta\bm{e}_{1}+\sin\zeta\bm{e}_{2})$. Here $\bm{e}_{1},\bm{e}_{2}$ are unit vector fields chosen to ensure $(\bm{e}_{1},\bm{e}_{2},\bm{b})$ is a right-handed orthonormal frame. Then set $\Sigma=\{(\bm{x},v_{\parallel},v_{\perp},0)\}$. We will refer to $\Sigma$ as the guiding center Poincaré section.
Since $\mathcal{R}|_{\epsilon=0}=R_{0}=-\partial_{\zeta}$, (I) $\Rightarrow$ (II) when $\epsilon=0$. Due to persistence of transversal intersections under deformations GP_1974 , (I) $\Rightarrow$ (II) for $\epsilon$ in some open interval $(-\epsilon_{0},\epsilon_{0})$, $\epsilon_{0}>0$, and $v_{\perp}$ not too close to 00. However, (II) is still essential because we cannot predict $\epsilon_{0}$ in advance.
@@
Intrinsically, a particle’s guiding center is equal to the $U(1)$-orbit passing through that particle’s phase space location. Guiding center phase space is therefore intrinsically the collection of all $U(1)$-orbits. Constructing a guiding center model, perturbative or non-perturbative, requires equipping this space with coordinates and identifying an evolution law for $U(1)$-orbits in that coordinate system. Traditional guiding center theory painstakingly constructs such coordinates order-by-order in $\epsilon$, resulting in non-unique model equations with exploding complexity that cannot be applied in the non-perturbative regime identified in Fig. 2.
@@
Breaking with tradition, we coordinatize $U(1)$-orbits by assigning to each orbit its point of intersection with $\Sigma$, $(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r,0)$. In this manner we regard $\Sigma$ as a concrete realization of the abstract space of $U(1)$-orbits. We will use $\pi$ to denote the map that sends each particle-space point $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ to its $U(1)$-orbit-mate on $\Sigma$, $\pi(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r)$. We will refer to $\pi$ as the footpoint map and the image of $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ under $\pi$ as that phase point’s footpoint. Note that $\pi(\bm{x},v_{\parallel},v_{\perp},0)=(\bm{x},v_{\parallel},v_{\perp})$, indicating that a particle’s footpoint coincides with its phase space location whenever the particle’s trajectory intersects $\Sigma$. We define the guiding center of any particle as that particle’s footpoint on $\Sigma$.
@@
Figure 3: (a,d) Poincaré sections of the full-orbit dynamics (black) and level sets of the learned $\mathcal{J}_{\Sigma}$ (color) for $\epsilon\in\{0.15,0.27\}$ (cf. Fig. 2 (b,c)). (b,e) Error of $\mathcal{J}_{\Sigma}$ prediction of the dynamics from the full-orbit dynamics over a gyroperiod. (c,f) Error of the order-$j$ truncated dynamics (4) over a gyroperiod with $j=(1,0)$.
@@
As a particle moves through phase space its footpoint traces an image curve on $\Sigma$. Because the Lorentz force equations of motion are $U(1)$-invariant by hypothesis that image curve is a streamline for a uniquely defined vector field on $\Sigma$, $X_{\Sigma}=\dot{\bm{X}}\cdot\partial_{\bm{X}}+\dot{u}\,\partial_{u}+\dot{r}\,\partial_{r}$. See Theorem 1 in supplementary material. We refer to $X_{\Sigma}$ as the footpoint flow field. By identifying $\Sigma$ with the space of $U(1)$-orbits we identify footpoint dynamics with $U(1)$-orbit dynamics as a byproduct. Thus, the components of the footpoint flow field define the non-perturbative guiding center equations of motion. We identify explicit formulas for these components involving only $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$ as follows.
@@
Any pair of functions $f,g$ on $\Sigma$ determines a pair of $U(1)$-invariant functions $F=f\circ\pi$, $G=g\circ\pi$ on all of phase space by precomposing with the footpoint map. Because $\mathcal{R}$ is a Hamiltonian vector field the function $\{F,G\}$ is also $U(1)$-invariant. It follows that the restriction $\{f,g\}_{\Sigma}=\{F,G\}\mid\Sigma$ defines a Poisson bracket on $\Sigma$. This bracket, together with the r estricted kinetic energy, $E_{\Sigma}=E\mid\Sigma=u^{2}/2+r^{2}/2$, determines a Hamiltonian system on $\Sigma$. Theorem 2 in supplementary material shows that this system coincides with the footpoint flow field. Thus, the components of the footpoint flow field $X_{\Sigma}$ may be computed explicitly given an explicit formula for the Poisson bracket $\{f,g\}_{\Sigma}$. Remarkably, this bracket can be expressed entirely in terms of $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$. The general formula is contained in the proof of Theorem 2 in supplementary material. For the $4D$ $(x,y,v_{\perp},\zeta)$-phase space appropriate for magnetic fields of the form $\bm{B}=B(x,y)\,\bm{e}_{z}$ the result is
@@
Here $\mathcal{J}_{\Sigma}=\mathcal{J}\mid\Sigma$ denotes restriction of the non-perturbative adiabatic invariant to the guiding center Poincaré section. We remark that these non-perturbative guiding center evolution laws are exact granted existence of the non-perturbative $U(1)$-symmetry generated by $\mathcal{R}$.
Although the formulas (2)-(3) reveal the central role played by the adiabatic invariant in guiding center modeling, they cannot be numerically simulated without an expression for $\mathcal{J}_{\Sigma}$. In the perturbative regime, $\epsilon\ll 1$, truncations of the magnetic moment asymptotic series, such as the second-order result Burby_gc_2013 for $\bm{B}=B(x,y)\,\bm{e}_{z}$,
@@
can be used to overcome this challenge. On the other hand, in the non-perturbative regime identified in Fig. 2, the traditional truncated series representation for $\mathcal{J}_{\Sigma}$ fails.
We instead choose to learn $\mathcal{J}_{\Sigma}$ directly from trajectories of the Lorentz force equations for (1). To do this, we minimize the Rayleigh quotient
@@
where $\mathcal{J}_{\Sigma}$ is discretized by $39\times 39\times 2$ Fourier by Fourier by Chebyshev modes in $\bm{r}=(x,y,r)\in\Omega=\Omega_{xy}\times[\sqrt{2}-0.01,\sqrt{2}+0.01]$ and $\Omega_{xy}$ is a periodic cell of the magnetic field. The first residual
@@
attempts to minimize the difference between the dynamics in $\mathcal{J}_{\Sigma}$ and and the full trajectory on the Poincaré section. A similar objective function appeared previously in Messenger_2024 , where a parametric averaged Hamiltonian appeared in place of our parametric $\mathcal{J}_{\Sigma}$. For the sum, $1000$ points $\hat{\bm{r}}_{i}$ are initially sampled from a low-discrepancy sequence on $\Omega$. Then, for the $\hat{N}$ points $\hat{\bm{r}}_{i}$ that correspond to integrable trajectories, footpoint time derivatives $(\dot{X}_{i},\dot{Y}_{i})$ are estimated using a technique based on Ruth:2024 , while the non-integrable trajectories are discarded (see supplementary materials). To define the second residual, let $F:\Omega\to\Omega$ be the Poincaré map from intersections of the Lorentz dynamics with $\Sigma$. We define $R_{\mathrm{Inv}}^{2}=\frac{1}{N}\sum_{i}(\mathcal{J}_{\Sigma}(\bm{r}_{i})-\mathcal{J}_{\Sigma}(F(\bm{r}_{i})))^{2}$ and $\lVert\mathcal{J}_{\Sigma}\rVert_{L^{2}(\Omega)}^{2}=\frac{1}{N}\sum_{i}\mathcal{J}_{\Sigma}(\bm{r}_{i})^{2}$, where the $\bm{r}_{i}$ are sampled on a $79\times 79\times 3$ Fourier by Fourier by Chebyshev-Lobatto quadrature grid on $\Omega$ and $N=51^{2}\times 9$. The residual $R_{\mathrm{Inv}}$ is minimized when $\mathcal{J}_{\Sigma}$ is invariant Ruth:2023 . This, along with the oversampling in $\bm{r}_{i}$, serves to smooth $\mathcal{J}_{\Sigma}$ in the chaotic regions. Note that both residuals in (5) are quadratic in $\mathcal{J}_{\Sigma}$, so the quotient can be minimized via a single generalized eigenvalue problem.
@@
Fig. 3 compares predictions using our learned $\mathcal{J}_{\Sigma}$ to ground-truth full-orbit simulation data previously shown in Fig. 2 at $\epsilon\in\{0.15,0.27\}$. The global phase portraits in Fig. 3 (a,d) reveal visually-obvious improvements over the invariants of the asymptotic theory shown in Fig. 2 (b,c). Note the elimination of unphysical protrusions present in level sets of the perturbative adiabatic invariant visible in Fig. 2 (b,c). To quantitatively measure the error in the invariant, let $F_{\mathcal{J}}$ be the map obtained by evolving Eqs. (2)-(3) over a gyroperiod for the learned invariant, and $F_{j}$ be the equivalent map for the order-$j$ truncation of (4). The log-absolute error $\log_{10}|F_{\mathcal{J}}-F|$ is plotted in Fig. 3 (b,e), which can be compared to the log-absolute error $\log_{10}|F_{j}-F|$ with $j=1$ for $\epsilon=0.15$ and $j=0$ for $\epsilon=0.05$ (higher orders result in zero denominators $D$ in both cases). For both values of $\epsilon$, the guiding center dynamics predicted using the learned $\mathcal{J}_{\Sigma}$ outperforms the traditional asymptotic theory by orders of magnitude over most of the phase portrait.
Our non-perturbative guiding center model offers transformative improvements in the optimization of stellarators for $\alpha$-particle confinement. It can improve trajectory accuracy in efforts to improve $\alpha$-confinement through direct calculation of $\alpha$-particle dynamics albert_accelerated_2020 ; albert_alpha_2023 . It promises to broaden the scope of advanced confinement concepts like quasisymmetry to include $\alpha$-particles, for instance by targeting symmetries of the non-perturbative guiding center model. We aim to realize these improvements by extending our $\mathcal{J}_{\Sigma}$-learning technique to general candidate stellarator configurations and by developing a data-driven functional representation of the non-perturbative adiabatic invariant, either by neural networks LuLu_2021 or sparse regression Brunton_2016 ; Messenger_2021 .

### v2

# Non-perturbative guiding center model for magnetized plasmas

@@ Abstract
Successful design of a stellarator reactor requires optimizing fusion-born $\alpha$-particle confinement to sustain fusion burn and minimize damage to plasma facing components. Stellarator design efforts therefore require an efficient model for $\alpha$-particle dynamics. State-of-the-art stellarator optimization routines employ the traditional guiding center model Kruskal_1962 ; Littlejohn_1981 ; Littlejohn_1983 ; Littlejohn_1984 ; Cary_2009 ; Burby_loops_JMP , which eliminates the short cyclotron timescale and provides the theoretical basis for advanced confinement concepts such as quasisymmetry Bo2_1983 ; NZ_1988 ; Burby_phase_2013 ; Helander_review_2014 ; BKM_2020 ; Rodriguez_2020 , omnigeneity Hall_1975 ; Helander_review_2014 ; Landreman_2012 ; Parra_2015 , quasi-isodynamicity QI_2023 , and isodrasticity BMN_2023 . This approximation assumes $\epsilon=\rho_{0}/L_{0}\ll 1$, where $\rho_{0}$ denotes a characteristic particle gyroradius and $L_{0}$ denotes a characteristic length scale for the magnetic field $\bm{B}$. When $\epsilon$ is not sufficiently small, traditional guiding center predictions may deviate significantly from those of full-Lorentz-force modeling, leading to imprecise estimates of particle confinement. Because the gyroradius scales linearly with particle speed, energetic $\alpha$-particles may fall in this problematic range even when thermal particles do not. Accuracy of the traditional guiding center model should be critically evaluated in any future stellarator design study.
Figure 1: (Top) Poincaré sections ($z=v_{z}=v_{y}=0,v_{x}>0$) of the full-orbit dynamics of Eq. (1) (black) and level sets of an adiabatic invariant truncation (color) initialized with $v_{x}=\epsilon$ for different values of $\epsilon$. Chaos appears in lobes with smaller field strength due to larger local-$\epsilon$. Lowest-order truncation is used in (a-c), consistent with guiding center models used in practical calculations. $y_{0}$ denotes $y$-coordinate of blue cross. (Bottom) Relative error time traces of three truncations of adiabatic invariant series for the same $\epsilon$ values. Blue crosses denote trajectories underlying (d-f).
Recent numerical studies Assuncao_2023 ; Rodrigues_2024 ; Carbajal_2024  [^1] Ref. Carbajal_2024 is available upon request. Contact: L. Carbajal, leo.carbajal@typeoneenergy.com. find significant deviations in computed $\alpha$-particle loss fractions in various reactor-scale stellarator configurations when using full-orbit vs traditional guiding center particle pushers. These and other calculations indicate that, depending on the stellarator configuration under study, guiding center predictions can either overestimate or underestimate the energetic particle loss fraction computed from full-orbit computations. While some configurations show no discrepancies at all, many lead to differences greater than $60\%$ of the guiding center prediction. Quasisymmetric designs tend to show greater discrepancies than quasi-isodynamic designs. Single trajectory comparisons Assuncao_2023 reveal particles with velocity vectors nearly perpendicular to the magnetic field account for the bulk of these discrepancies. Current designs of some optimized stellarator fusion power plants Hegna_2025 ; Guttenfelder_2025 show good confinement of fusion-born alpha-particles with less than few percent of alpha-particle energy losses. Nevertheless, discrepancy between full-orbit and guiding-center estimates of alpha-particle confinement is still observed Carbajal_2025 , suggesting improvements over guiding center are still necessary for high-fidelity evaluations.
Such deviations may appear surprising because estimating $\epsilon$ for $\alpha$-particles in existing stellarator reactor concepts using nominal minor radius as magnetic scale length $L_{0}$ leads to $\epsilon\sim 2\text{-}3\%$. However, modern stellarator designs often feature highly elongated and curved poloidal cross sections. The local $\epsilon$ in those regions may exceed the nominal value by a factor of  ${\sim}\,5$, leading to a bleaker view of guiding center validity. For example, detailed phase portraits of $\alpha$-particle trajectories in the $2D$ magnetic field $\bm{B}=B(x,y)\,\bm{e}_{z}$, with
@@
$k_{x1}=k_{y2}=3,\quad k_{y1}=k_{x2}=1,\quad a_{1}=a_{2}=0.3,$
reveal significant shifts in particle drift surfaces away from traditional guiding center predictions for $\epsilon\sim 15\%$ (note that $B$ is scaled so $L_{0}=1$). See Fig. 1. Since bounce motion and field line curvature play no role in this field, the quality of traditional guiding center predictions at $\epsilon\sim 15\%$ should only degrade further in a fully $3D$ configuration. Worse still, Fig. 1 reveals that including higher-order perturbative corrections Burby_gc_2013 to the traditional guiding center model, which represents a low-order truncation, offers limited accuracy improvements, if any. This signifies collapse of the optimal truncation order Berry_1991 for the adiabatic invariant asymptotic series. These observations motivate development of a non-perturbative variant of guiding center theory to replace traditional guiding center when it fails.
@@
This Letter describes a non-perturabative guiding center model suitable for $\alpha$-particles in stellarators, runaway electrons in tokamaks Carbajal_2017 ; Carbajal_2020 , and magnetically-confined energetic particles Ogawa_2016 produced by neutral beams Pefferle_2015 , ion cyclotron resonance heating, or fusion reactions Cecconello_2018 more broadly. First we deduce formally-exact non-perturbative guiding center equations of motion assuming a hidden symmetry with associated conserved quantity $\mathcal{J}$. We refer to $\mathcal{J}$ as the non-perturbative adiabatic invariant. These equations are completely determined by first-order derivatives of $\mathcal{J}$ and the magnetic field. They enjoy Hamiltonian structure comparable with that of asymptotic theory Littlejohn_1981 . Then we describe a data-driven method for learning $\mathcal{J}$ from a dataset of full-orbit $\alpha$-particle trajectories. We apply this method to the $\alpha$-particle dynamics shown in Fig. 1 and find the learned non-perturbative guiding center model significantly outperforms the standard guiding center expansion. Our proposed method for learning $\mathcal{J}$ applies on a per-magnetic field basis; changing $\bm{B}$ requires re-training. This makes it well-suited to stellarator design assessment tasks, such as $\alpha$-loss-fraction uncertainty quantification.
@@
In a seminal paper Kruskal_1962 , M. Kruskal showed traditional guiding center theory originates from a hidden perturbative $U(1)$-symmetry of the ordinary differential equations describing single-particle motion, $X=\dot{\bm{x}}\cdot\partial_{\bm{x}}+\dot{\bm{v}}\cdot\partial_{\bm{v}}$, where $\dot{\bm{v}}=\bm{v}\times\bm{B}$ and $\dot{\bm{x}}=\epsilon\,\bm{v}$. Here $U(1)$-symmetry refers to a periodic one-parameter group of phase space transformations that maps solutions of the Lorentz force equations into other solutions. At leading order, the $U(1)$-symmetry is concretely expressed via gyrophase translations. Kruskal dubbed the infinitesimal generator of the perturbative hidden symmetry the roto rate $R$ and showed that its formal expansion in powers of $\epsilon$ is uniquely determined to all orders, with first term given by $R_{0}=\bm{v}\times\bm{b}\cdot\partial_{\bm{v}}$.
@@
i.e. $\{E,\mathcal{J}\}=0$. We refer to $\mathcal{J}$ as the non-perturbative adiabatic invariant.
@@
To ensure topological similarity with the known $U(1)$-symmetry at $\epsilon=0$, we also assume (II) each $U(1)$-orbit generated by $\mathcal{R}$ uniquely and transversally intersects a certain phase space Poincaré section $\Sigma$ defined as follows. Introduce phase space coordinates $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ such that $\bm{v}=v_{\parallel}\,\bm{b}+v_{\perp}\,(\cos\zeta\bm{e}_{1}+\sin\zeta\bm{e}_{2})$. Here $\bm{e}_{1},\bm{e}_{2}$ are unit vector fields chosen to ensure $(\bm{e}_{1},\bm{e}_{2},\bm{b})$ is a right-handed orthonormal frame. Then set $\Sigma=\{(\bm{x},v_{\parallel},v_{\perp},0)\}$. We refer to $\Sigma$ as the guiding center Poincaré section.
@@
Intrinsically, a particle’s guiding center is equal to the $U(1)$-orbit passing through that particle’s phase space location. Guiding center phase space is therefore intrinsically the collection of all $U(1)$-orbits. Constructing a guiding center model, perturbative or non-perturbative, requires equipping this space with coordinates and identifying an evolution law for $U(1)$-orbits in that coordinate system. Traditional guiding center theory painstakingly constructs such coordinates order-by-order in $\epsilon$, resulting in non-unique model equations Parra_Calvo_2014 with exploding complexity that cannot be applied in the non-perturbative regime identified in Fig. 1.
@@
Breaking with tradition, we label $U(1)$-orbits by assigning to each orbit its point of intersection with $\Sigma$, $(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r,0)$, or footpoint. In this manner we regard $\Sigma$ as a concrete realization of the abstract space of $U(1)$-orbits. A particle’s footpoint is just the footpoint of the $U(1)$-orbit passing through that particle’s phase space location. We denote the footpoint of a particle at $(\bm{x},v_{\parallel},v_{\perp},\zeta)$ as $\pi(\bm{x},v_{\parallel},v_{\perp},\zeta)=(\bm{X},u,r)$. Note that $\pi(\bm{x},v_{\parallel},v_{\perp},0)=(\bm{x},v_{\parallel},v_{\perp})$, indicating that a particle’s footpoint coincides with its phase space location whenever the particle’s trajectory intersects $\Sigma$, which happens once each gyroperiod. We define the guiding center of any particle as that particle’s footpoint on $\Sigma$.
@@
Figure 2: (a,d) Poincaré sections of the full-orbit dynamics (black) and level sets of the learned $\mathcal{J}_{\Sigma}$ (color) for $\epsilon\in\{0.15,0.27\}$ (cf. Fig. 1 (b,c)). (b,e) Error of $\mathcal{J}_{\Sigma}$ prediction of the dynamics from the full-orbit dynamics over a gyroperiod. (c,f) Error of the order-$j$ truncated dynamics (4) over a gyroperiod with $j=(1,0)$. The error is computed on a uniform $201\times 201$ point grid — a higher resolution than the training data.
@@
As a particle moves through phase space its footpoint traces a curve on $\Sigma$. Because the Lorentz force equations of motion are $U(1)$-invariant by hypothesis that curve is a streamline for a uniquely defined vector field on $\Sigma$ that we call the footpoint flow field, $X_{\Sigma}=\dot{\bm{X}}\cdot\partial_{\bm{X}}+\dot{u}\,\partial_{u}+\dot{r}\,\partial_{r}$. For a proof, see Theorem 1 in supplementary material. The components of the footpoint flow field define our non-perturbative guiding center equations of motion. We identify explicit formulas for these components involving only $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$ as follows. See Figure 1 in the supplementary material for an illustration of the phase space geometry of footpoint flow.
@@
Any pair of functions $f,g$ on $\Sigma$ determines a pair of $U(1)$-invariant functions $F=f\circ\pi$, $G=g\circ\pi$ on all of phase space by precomposing with the footpoint map $\pi$. Because $\mathcal{R}$ is a Hamiltonian vector field the function $\{F,G\}$ is also $U(1)$-invariant. It follows that the restriction $\{f,g\}_{\Sigma}=\{F,G\}\mid\Sigma$ defines a Poisson bracket on $\Sigma$. This bracket, together with the restricted kinetic energy, $E_{\Sigma}=E\mid\Sigma=u^{2}/2+r^{2}/2$, determines a Hamiltonian system on $\Sigma$. Theorem 2 in supplementary material shows this system coincides with the footpoint flow field. Thus, components of the footpoint flow field $X_{\Sigma}$ may be computed explicitly given an explicit formula for the Poisson bracket $\{f,g\}_{\Sigma}$. This bracket can be expressed entirely in terms of $\mathcal{J}$, $\bm{B}$, and the unit vectors $\bm{e}_{1},\bm{e}_{2}$. The general formula is contained in the proof of Theorem 2 in supplementary material, which includes Ref. Guillemin_Sternberg_1980 . For the $4D$ $(x,y,v_{\perp},\zeta)$-phase space appropriate for magnetic fields of the form $\bm{B}=B(x,y)\,\bm{e}_{z}$ the result is
@@
Here $\mathcal{J}_{\Sigma}=\mathcal{J}\mid\Sigma$ denotes restriction of the non-perturbative adiabatic invariant to the guiding center Poincaré section. These non-perturbative guiding center evolution laws are exact granted existence of the non-perturbative $U(1)$-symmetry generated by $\mathcal{R}$.
Although formulas (2)-(3) reveal the central role played by the adiabatic invariant in guiding center modeling, they cannot be numerically simulated without an expression for $\mathcal{J}_{\Sigma}$. In the perturbative regime, $\epsilon\ll 1$, truncations of the magnetic moment asymptotic series, such as the second-order result Burby_gc_2013 for $\bm{B}=B(x,y)\,\bm{e}_{z}$,
@@
provide this missing ingredient. On the other hand, Fig. 1 demonstrates inadequacy of truncated series representations in the non-perturbative regime.
We instead choose to learn $\mathcal{J}_{\Sigma}$ directly from trajectories of the Lorentz force equations for (1) by minimizing the Rayleigh quotient
@@
where $\mathcal{J}_{\Sigma}$ is discretized by $39\times 39\times 2$ Fourier by Fourier by Chebyshev modes in $\bm{r}=(x,y,r)\in\Omega=\Omega_{xy}\times[\sqrt{2}-0.01,\sqrt{2}+0.01]$ and $\Omega_{xy}$ is a periodic cell of the magnetic field. Here $\lVert\mathcal{J}_{\Sigma}\rVert_{L^{2}(\Omega)}^{2}=\frac{1}{N}\sum_{i}\mathcal{J}_{\Sigma}(\bm{r}_{i})^{2}$, where the $\bm{r}_{i}$ are sampled from a $79\times 79\times 3$ Fourier by Fourier by Chebyshev-Lobatto quadrature grid on $\Omega$ and $N=51^{2}\times 9$. The first residual
@@
penalizes differences between predictions from $\mathcal{J}_{\Sigma}$ and full-orbit trajectories on the Poincaré section. A similar objective function appeared previously in Messenger_2024 , where a parametric averaged Hamiltonian appeared in place of our parametric $\mathcal{J}_{\Sigma}$. For the sum we initially sample $1000$ points $\hat{\bm{r}}_{i}$ from a low-discrepancy sequence on $\Omega$. For the $\hat{N}$ points that correspond to integrable trajectories, we estimate footpoint time derivatives $(\dot{X}_{i},\dot{Y}_{i})$ using a technique based on Ruth:2024 . (See supplementary materials.) The residual $R_{\mathrm{Dyn}}^{2}$ ignores non-integrable trajectories. The second residual is given by $R_{\mathrm{Inv}}^{2}=\frac{1}{N}\sum_{i}(\mathcal{J}_{\Sigma}(\bm{r}_{i})-\mathcal{J}_{\Sigma}(F(\bm{r}_{i})))^{2}$, where $F:\Omega\to\Omega$ denotes the full-orbit Poincaré map on $\Sigma$. It penalizes non-invariance of $\mathcal{J}_{\Sigma}$ Ruth:2023 . This, along with the oversampling in $\bm{r}_{i}$, smoothes $\mathcal{J}_{\Sigma}$ in chaotic regions. The quotient can be minimized via a single generalized eigenvalue problem because both residuals in (5) are quadratic in $\mathcal{J}_{\Sigma}$. Computing the $\hat{N}$ samples of $({X},{Y})$ dominates computational time. An average trajectory computation requires around $1000$ gyroperiods, so the method’s total cost is on the order of the time to simulate $1000$ gyroperiods for $1000$ particles using a full-orbit integrator.
Many other methods compute approximate invariant structures, including Lagrangian coherent structures hadjighasem2017 ; froyland2014 , flux surfaces dewar-qfm-maps ; dewar-qfm-fields ; Ruth:2023 , and invariant tori more generally haro-parameterization . However, these methods do not distinguish between the true adiabatic invariant $\mathcal{J}$ associated with continuation of Kruskal’s perturbative $U(1)$-symmetry and arbitrary invariants of the form $I(E,\mathcal{J})$. Our method accounts for this ambiguity because replacing $\mathcal{J}$ with $I(E,\mathcal{J})$ in the non-perturbative equations of motion (2) generally changes predicted footpoint trajectories, and thereby increases the residual $R_{\mathrm{Dyn}}^{2}(\mathcal{J}_{\Sigma})$. In contrast to general constants of motion, the learned non-perturbative invariant $\mathcal{J}$ completely determines the dynamics.
@@
Fig. 2 compares predictions using learned $\mathcal{J}_{\Sigma}$ to full-orbit simulation data previously shown in Fig. 1 at $\epsilon\in\{0.15,0.27\}$. The global phase portraits in Fig. 2 (a,d) reveal significant improvements over the asymptotics shown in Fig. 1 (b,c). Let $F_{\mathcal{J}}$ denote the approximate Poincaré map obtained by evolving Eqs. (2)-(3) over a gyroperiod using learned $\mathcal{J}$. Let $F_{j}$ denote the corresponding map using order-$j$ truncation of (4). Figure 2 (b,e) compares the log-absolute errors, $\log_{10}|F_{\mathcal{J}}-F|$ and $\log_{10}|F_{j}-F|$, with $j=1$ for $\epsilon=0.15$ and $j=0$ for $\epsilon=0.27$ (larger $j$ leads to zero denominators $D$ in both cases). For both $\epsilon$-values, Poincaré map dynamics predicted using learned $\mathcal{J}_{\Sigma}$ outperforms traditional asymptotics by orders of magnitude over most of the phase portrait.
Discussion– Advanced stellarator design principles like quasisymmetry and omnigeneity aim to improve confinement by imbuing the traditional guiding center model with desirable dynamical features, such as symmetry or small radial flux. Reliance on traditional guiding center limits applicability of these principles to the regime $\epsilon\ll 1$. New confinement principles targeting comparable dynamical improvements for the non-perturbative guiding center model promise to extend the benefits of optimized confinement to larger-$\epsilon$ particles, such as $\alpha$-particles in stellarators that break the traditional guiding center model.
Assessing a candidate stellarator reactor design merits sensitivity analysis of $\alpha$-particle loss fraction over an energy confinement time. Assessing sensitivity to perturbations of the $\alpha$-particle distribution function requires either high-dimensional adjoint solves or large Monte-Carlo (MC) sampling of $\alpha$-particle trajectories from various perturbed distributions. Because each MC sample evolves in the same field configuration, computational overhead associated with training the non-perturbative guiding center model falls well below the cost of performing all required Monte-Carlo sampling using a full-orbit approach. Thus, in device designs where guiding center fails for $\alpha$-particles, accuracy and timestepping efficiency motivate trajectory sampling using the non-perturbative guiding center model over traditional guiding center or full-orbit models.
Direct optimization of $\alpha$-particle confinement albert_accelerated_2020 ; albert_alpha_2023 ; bindel_direct_2023 entails repeated ensemble trajectory simulations during an inner field-optimization loop. Computational feasibility of this technique currently requires traditional guiding center for trajectory calculations, limiting its applicability to particles with $\epsilon\ll 1$. Deploying a trained non-perturbative guiding center model in place of traditional guiding center promises to extend applicability of direct fast-particle confinement optimization to particles with non-perturbative $\epsilon$. However, the training method proposed here falls short for inner-loop applications because it must be repeated whenever the field changes, thereby encurring unacceptable computational overhead comparable to full-orbit trajectory generation. Overcoming this barrier requires solving the operator learning problem LuLu_2021 of mapping $\bm{B}$ to $\mathcal{J}$. Recent advances in machine learning, especially in foundation models Subramanian_2023 and sparse regression Brunton_2016 ; Messenger_2021 , stand poised to tackle this challenge.
[…]